Typical blood velocities in the coronary arteries
Answer #1

a.

F = IqI v B

where,

q = 1.6 x 10^-19 (this is the charge of proton )

v = 0.15 m/s

B = 0.22 T

F = 1.6 x 10^-19 x 0.22 x 0.15

F = 5.28 x 10^21 N

b.

F = qE

E = F/q

E = (5.28 x 10^-21)/ (1.6*10^-19)

E = 0.033 V/m = 3.3 x 10^-2 N/C

c.

3.3 × 10−2 V/m x 0.0034 m = 1.112 x 10^-4 V